# StereoMath: An Accessible and Musical Equation Editor

### v1

@@ Abstract.
†This work was initiated prior to enrollment and conducted independent of Carnegie Mellon, and conducted without support from the institution
@@ 2. Related Work
Most related work in this space relies on LaTeX coding. For example, ALAP (Manzoor et al., 2019) focuses on improving the LaTeX debugging experience, and tools such as AROW (Seo and McCurry, 2019) and WriteR (Godfrey, 2024) offer easier document layout editing, but still rely on LaTeX coding for equations. However, LaTeX is known to have a steep learning curve, with extra challenges for BLV users (Seo and McCurry, 2019).
@@
Some tools, such as JAWS MathViewer and ChromeVox-NavMatBr (da Paixão Silva et al., 2018), organize mathematical expressions into a tree, allowing faster and more comprehensible navigation (da Paixao Silva et al., 2017). Unfortunately, JAWS does not adhere to the MathSpeak protocol, resulting in ambiguous readalouds: for example, “$\sqrt{3i}-2i$” and “$\sqrt{3i-2i}$” result in the same output (noa, 2024b, c).
Luckily, work in other domains has successfully overcome similar challenges, using sonification (Holloway et al., 2022; Barrass and Kramer, 1999) (for example, in SAS Graphics Accelerator (noa, 2024e) and EmacsSpeak (Raman, 1996)), as well as employing modality switching (Paivio, 1991), which has been shown to reduce cognitive load (Cao et al., 2009; Klatzky et al., 2006; Brünken et al., 2004).
To the best of our knowledge, this work is the first equation editor built to mirror the intuitive nature of sighted equivalents, in a 2D spatial sense. The only web-based alternative, Desmos (noa, 2024a), is closed-source and lacks LaTeX code generation.
@@ 4.1. Overview
[…]
[A diagram detailing the different elements of the equation editor]A screenshot of the equation editor, along with arrows and boxes highlighting different elements. Elements are as follows: A box highlighting a node, with caption ”tree-based representation, with features such as loopback buffer”; a box highlighting an input field with caption ”custom JS input field allows for application-specific behavior and custom spatial readalouds”; a box highlighting the vertical parentheses bars with caption ”Every mathematical element has corresponding “grouping symbols.” Audio tone plays in corresponding spatial location”; a box highlighting the spatial cursor, with caption ”Spatial cursor, used for keyboard-based spatial navigation. Tones represent position in equation table”. Finally, the image also highlights the Settings and LaTeX generation buttons, as well as the live aria messages
@@ 4.2. Architecture
[Architecture overview, with trees and spatial info]A diagram featuring the internal equation tree on the left side, and the spatial information on the right side, which is connected to a custom TTS on the server-side. The internal nodes on the equation tree are labeled as operators, and the leaf nodes are input fields. It also shows an example custom input field, and an example series of musical chords representing an earcon.
On the server side, our biggest architectural hurdle was implementing our custom TTS. We piped announcements to the server, generated raw PCA audio on the server using Piper TTS (noa, 2024d), and sent it back to the client. The client would then use the Web Audio API and Tone.js (noa, 2024f) to make the binaural and pitch adjustments, respectively.
Custom Input Fields We developed a custom input field and overrode the built-in screen reader feedback. Although normally discouraged, this was necessary for a few reasons.
First, we needed custom hooks to determine exactly when the cursor moved, to override the default behavior upon falling off either edge of the input field. This let us seamlessly navigate across our tree-based space. The built-in HTML input element was much less consistent and added extraneous and confusing announcements such as ”end of text.”
[…]
Third, and perhaps most importantly, it allowed us to modify and customize the speech output audio, which was critical for spatial sound. For example, consider an input field with the following text, with the ”—” representing the caret: ”a—bcde”. If the user were to scan character by character to the end, the user would notice that the audio moves from left to right, with ”e” being played fully on the right speaker. Thus, we hijack every possible access point and make the entire user experience spatial.
@@
Keyboard Spatial Navigation Algorithm This provides a muscle memory, tactile interface while reusing the user’s own keyboard. We take the middle four rows of the keyboard, and, based on the user’s keyboard layout, find the physical relative position of each key. A high level example is as follows: Take the equation $\frac{a}{b}+\frac{c}{d}+\frac{e}{f}$. Notice that it is essentially a grid, with two rows and three columns. If the user presses a key in the top right of the keyboard, such as ”p,” the input field containing ”$e$” will be highlighted. In addition, depending on whether the user is in ”row mode” or ”column mode,” the editor will also play a series of notes representing the number of elements in the current row/column, as well as another note to indicate which item is currently selected.

### v2

@@ 2. Related Work
Most related works in this space rely on LaTeX coding. For example, ALAP (Manzoor et al., 2019) focuses on improving the LaTeX debugging experience, and tools like AROW (Seo and McCurry, 2019) and WriteR (Godfrey, 2024) offer easier document layout editing, but still rely on LaTeX coding for equations. However, LaTeX is known to have a steep learning curve, with extra challenges for BLV users (Seo and McCurry, 2019).
@@
Some tools, such as JAWS MathViewer and ChromeVox-NavMatBr (da Paixão Silva et al., 2018), organize mathematical expressions into a tree, enabling faster and more comprehensible navigation (da Paixao Silva et al., 2017).Unfortunately, JAWS does not adhere to the MathSpeak protocol, resulting in ambiguous readalouds: e.g., $\sqrt{3i}-2i$” and $\sqrt{3i-2i}$” result in the same output (14; 15).
Luckily, work in other domains has successfully overcome similar challenges, by using sonification (Holloway et al., 2022; Barrass and Kramer, 1999) (e.g., in SAS Graphics Accelerator (29) and EmacsSpeak (Raman, 1996)), as well as employing modality switching (Paivio, 1991), which has been shown to reduce cognitive load (Cao et al., 2009; Klatzky et al., 2006; Brünken et al., 2004).
To our best knowledge, this work is the first equation editor built to mirror the intuitive nature of sighted equivalents. The only web-based alternative, Desmos (10), is closed-source and lacks LaTeX code generation.
@@ 4.1. Overview
Refer to caption A diagram detailing the different elements of the equation editor A screenshot of the equation editor, along with arrows and boxes highlighting different elements. Elements are as follows: A box highlighting a node, with caption ”tree-based representation, with features such as loopback buffer”; a box highlighting an input field with caption ”custom JS input field allows for application-specific behavior and custom spatial readalouds”; a box highlighting the vertical parentheses bars with caption ”Every mathematical element has corresponding “grouping symbols.” Audio tone plays in corresponding spatial location”; a box highlighting the spatial cursor, with caption ”Spatial cursor, used for keyboard-based spatial navigation. Tones represent position in equation table”. Finally, the image also highlights the Settings and LaTeX generation buttons, as well as the live aria messages
@@ 4.2. Architecture
Refer to caption Architecture overview, with trees and spatial info A diagram featuring the internal equation tree on the left side, and the spatial information on the right side, which is connected to a custom TTS on the server-side. The internal nodes on the equation tree are labeled as operators, and the leaf nodes are input fields. It also shows an example custom input field, and an example series of musical chords representing an earcon.
On the server-side, our biggest architectural hurdle was implementing our cusotm TTS. We piped announcements to the server, generated raw PCA audio on the server using Piper TTS (27), and sent it back to the client. The client would then use the Web Audio API and Tone.js (35) to make the binaural and pitch adjustments respectively.
Custom Input Fields We developed a custom input field and overrode the built-in screen reader feedback. Although normally discouraged, this was necessary for a few reasons:
First, we needed custom hooks to determine exactly when the cursor moved, to override the default behavior upon falling off either edge of the input field. This let us seamlessly navigate across our tree-based space. The built-in HTML input element was much less consistent, and added extraneous and confusing announcements such as ”end of text.”
@@
Third, and perhaps most importantly, it allowed us to modify and customize the speech output audio, which was critical for spatial sound. For example, consider an input field with the following text, with the ”—” representing the caret: ”a—bcde”. If the user were to scan, character by character, to the end, the user would notice that the audio moves from left to right, with ”e” being played fully on the right speaker. Thus, we hijack every possible access point, and make the entire user experience spatial.
@@
Keyboard Spatial Navigation Algorithm This provides a muscle memory, tactile interface while reusing the user’s own keyboard. We take the middle four rows of the keyboard, and, based on the user’s keyboard layout, find the physical relative position of each key. A high level example is as follows: Take the equation $\frac{a}{b}+\frac{c}{d}+\frac{e}{f}$. Notice that it is essentially a grid, with two rows and three columns. If the user presses a key in the top right of the keyboard, such as ”p,” the input field containing ”ee” will be highlighted. In addition, depending on whether the user is in ”row mode” or ”column mode,” the editor will also play a series of notes representing the number of elements in the current row/column, as well as another note to indicate which item is currently selected.